Ages Number of students
15-18 10
19-22 5
23-26 9
27-30 7
31-34 6
35-38 10


Find the relative frequency for the class with lower class limit 19
Relative Frequency =  %
Give your answer as a percent, rounded to two decimal places
Answer #1

N=10+5+9+7+6+10 =47

Relative frequency = Class frequency/Total frequency

Therefore,

The relative frequency for the class with lower class limit 19

=5/47

=0.1064

Relative Frequency = 10.64 %